6.407 g of solid CO2 (44.01 g/mol) is put in an empty sealed 3.45 x 103 mL container at a temperature of 305.7-K. When all the solid CO2 becomes gas, what will be the pressure (in bar) in the container? ( R = 0.08206 L.atm/mol.K)


1
Expert's answer
2021-06-18T07:33:02-0400

Given data in above 

Mass of CO(w) = 6.407 g 

Molar mass of CO(M)= 44.01 g/mol 

Volume of container (V) = 3.45×10mL

1 L = 1000 mL 

Then , 3.45×103 mL = (1L÷1000 mL)×3.45×10mL 

 = 3.45 L 

Temperature (K) = 305.7 K 

R ( Gas constant) = 0.08206 L .atm /mol. K 

Calculation of pressure of the COgas in container .

  We first calculate moles (n) of CO

"n =\frac {w}{M}"


 = "\frac{6.407 g}{44.01 g}"  

 = 0.15 mol 

  Now, using ideal gas equation to calculate pressure of gas in the container. 

PV= nRT 

P =  "\frac{nRT}{V }"

 "=\frac{0.15 mol×0.08206 L/ atm /mol/K×305.7 K) }{3.45 L }"


 = "\frac{3.76 atm}{3.45}"

 = 1.09 atm 

P = 1.09 atm 

We know that 

1 atm = 1.013 bar 

Then ,  "1.09 atm =\frac{1.013 bar× 1.09 atm }{1atm}"


        = 1.104 bar 

Hence, the pressure of the CO2 gas in the container is 1.104 bar .